OpenStudy (anonymous):

can anyone graph y = 2x^3 + 3x^2 – x + 2 fan and medal awarded !

OpenStudy (anonymous):

 almost like this

OpenStudy (whpalmer4):

Make yourself a table of x and y values. Have x go from -4 to 4, and the y value will be the value of the function when you plug in the value of x. x y -4 -3 -2 2(-2)^3+3(-2)^2 - (-2) + 2 = -16+12+2+2 = 0 -1 0 2(0)^3+3(0)^2-(0) + 2 = 2 1 2 2(2)^3+3(2)^2-(2) + 2 = 16+12-2+2 = 28 3 4 I've done a few points for you. When your table is complete, mark those points on your graph and draw a smooth curve through them. It will have a shape sort of like drawn above, but it doesn't actually cross the x-axis so many times.

OpenStudy (anonymous):

ive already told u @whpalmer4 i dont want your help. u didnt help AT ALL last time. and ive already gotten it done.

OpenStudy (anonymous):

and btw im still waiting for that answer. its been what ? 3 weeks now ? 4 ?

OpenStudy (anonymous):

i think u ow me something

OpenStudy (anonymous):

yes ma'am :)

OpenStudy (anonymous):

*owe

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

have a great day